# Chiral anomaly in inhomogeneous systems with nontrivial momentum space topology

### v2

@@ VI Conclusion and discussion
It would be interesting to extend this research in several directions. In particular, the role of interactions may be investigated. Their consideration is assumed to become the subject of a future work. Then the violation of chiral symmetry due to the chiral anomaly together with the absence of scale symmetry might cause the chiral symmetry breakdown at the level of the effective Dirac operator (that is inverse to the two point Green function with interaction corrections). At low energies the mass term will appear, i.e. the term with the transition between the left – handed and the right – handed fermions. If the non – Abelian gauge field entering expression for the chiral anomaly is dynamical, then this mass term is suppressed when the instanton density is suppressed. For the gauge field with the standard Yang – Mills action this occurs in the weak coupling regime. Then the chiral symmetry is still present on the level of effective action. We expect [18], that in this situation the result of the present paper should survive in some form.
Furthermore, it would be worthwhile to consider cases where the inhomogeneity cannot be removed through smooth deformations. We expect that in such situations, (81) may still hold in some form under less restrictive conditions than those considered above. However, a comprehensive analysis of the conditions required for the continued validity of (81) lies beyond the scope of the present paper. In this respect it is worth mentioning that the topological invariant of the form of (80) is known in the noncommutative geometry as pairing of the Connes - Chern character with a cyclic cocycle of the cyclic cohomology theory developed by Alain Connes [31, 32]. This theory may, presumably, be used for the calculation of the value of $N_{3}$ in rather general cases. More simple technique that may be applied to a certain class of such systems has been developed for such a calculation recently in [33].

### v3

@@ VI Conclusion and discussion
It would be interesting to extend this research in several directions. In particular, the role of interactions may be investigated. Their consideration is assumed to become the subject of a future work. Then the violation of chiral symmetry due to the chiral anomaly together with the absence of scale symmetry might cause the chiral symmetry breakdown at the level of the effective Dirac operator (that is inverse to the two point Green function with interaction corrections). At low energies the mass term will appear, i.e. the term with the transition between the left – handed and the right – handed fermions. If the non – Abelian gauge field entering expression for the chiral anomaly is dynamical, then this mass term is suppressed when the instanton density is suppressed. For the gauge field with the standard Yang – Mills action this occurs in the weak coupling regime. Then the chiral symmetry is still present on the level of effective action. We expect [18], that in this situation the result of the present paper should survive in some form (in the presence of additional non - dynamical external gauge field). In the opposite case of strong coupling Eq. (81) may be used, when we consider fermions in the presence of the given fixed configuration of the dynamical gauge field. (The divergence of axial current then becomes a dynamical variable as well as the gauge field.) Then Eq. (81) may even become useful in consideration of various non - perturbative effects, say, being unified with the technique of central or Abelian projection (see, for example, [31] and references therein).
Furthermore, it would be worthwhile to consider cases where the inhomogeneity cannot be removed through smooth deformations. We expect that in such situations, (81) may still hold in some form under less restrictive conditions than those considered above. However, a comprehensive analysis of the conditions required for the continued validity of (81) lies beyond the scope of the present paper. In this respect it is worth mentioning that the topological invariant of the form of (80) is known in the noncommutative geometry as pairing of the Connes - Chern character with a cyclic cocycle of the cyclic cohomology theory developed by Alain Connes [32, 33] (the other version of this formalism may be found in [9]). This theory may, presumably, be used for the calculation of the value of $N_{3}$ in rather general cases. More simple technique that may be applied to a certain class of such systems has been developed for such a calculation recently in [34].
@@
Finally, we would like to mention the potential possibility to extend our results to the case of the other currents, which are conserved classically. In particular, for the non - interacting systems with relativistic symmetry one can consider the currents generated by zilch fields [39, 40, 41, 42]. In case of the more complicated systems under certain conditions multiple classically conserved currents may be defined as well [43]. Some of the currents associated with zilch, as well as the other classically conserved currents may be anomalous [44, 45]. Correspondingly, the appearance of an analogue of our Eq. (81) might be expected. However, these (potentially anomalous) currents contain the algebraic structures that are not present in the ordinary axial current. Presumably, such structures are to replace $\gamma^{5}$ in Eq. (80). It is not clear whether the resulting expression in general case will remain a topological invariant. It could be, however, that it will remain robust to the smooth modifications of the system that respect symmetries resulted in the appearance of the corresponding classically conserved current. The consideration of this issue, however, is far out of the scope of the present paper.